OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now.

OpenStudy (anonymous):

WELCOME TO OS :D

OpenStudy (anonymous):

Ralph is... 18 I believe..

OpenStudy (anonymous):

Right now: Sara's age = x while Ralph's age = 3x In 6 years: Sara's age will be x + 6; and Ralph's age will be 3x + 6 So your equation is 3x + 6 = 2(x + 6) 3x + 6 = 2x + 12 x = 6 Ralphs age now is 3x or 3(6) = 18.